OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m=8, (7,1)

OpenStudy (whpalmer4):

Use the formula \(y-y_0 = m(x-x_0)\) to write the equation of a line with slope \(m\) passing through point \((x_0,y_0)\)

OpenStudy (anonymous):

I thought is was y-y1=m(x-x1) why did you put 0

OpenStudy (whpalmer4):

It's just a label. Exactly the same result...

OpenStudy (anonymous):

I need help getting the result

OpenStudy (whpalmer4):

What do you get if you plug in your numbers in your formula?

OpenStudy (anonymous):

I have no clue.

OpenStudy (whpalmer4):

You have the formula. You know that m = 8. What do you get if you rewrite the formula with m = 8?

OpenStudy (whpalmer4):

Wherever you see m, replace it with 8. What do you get?

OpenStudy (whpalmer4):

After that, replace x1 with 7, and y1 with 1 (those are the x and y values you are given for the point your line goes through).

OpenStudy (anonymous):

y-1=8x-56

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Probably rewrite that as \[y = 8x-55\]

OpenStudy (anonymous):

y=8x+55

OpenStudy (anonymous):

got my sign wrong

OpenStudy (whpalmer4):

no.... \[y-1=8x-56\]add 1 to each side \[y-1+1=8x-56+1\]\[y=8x-55\]

OpenStudy (whpalmer4):

Haste makes waste :-)

OpenStudy (anonymous):

I seem to have trouble with the signs cool thanks

OpenStudy (whpalmer4):

Self-awareness is good, use it to your advantage! If you know signs give you trouble, always check them.

OpenStudy (anonymous):

yeap